please help me with this problem....
12(x-4)-(x+2)=(x+2)(x-4)
solve for x using the quadratic formula.

Answer by rfer(16322)

12x-48-x-2=x^2-4x+2x-8
11x-50=x^2-2x-8
x^2-13x+42=0
Can you take it from here?